3. Simple versus compound interest Aa Aa Financial contracts involving investments, mortgages, loans, and so on are based on either a fixed or a variable interest rate. Assume that fixed interest rates are used throughout this question Olivia deposited $1,700 in a savings account at her bank. Her account will earn an annual simple interest rate of 7.8%. If she makes no additional deposits or withdrawals, how much money will she have in her account in 11 years? O $232.60 O $1,842.94 O $3,883.79 o $3,158.60 Now, assume that Olivias savings institution modifies the terms O$3,158.60 of her account and agrees to pay 7.8% in compound interest on $302.94 her $1,700 balance. All other things being equal, how much money will Olivia have in her account in 11 years? O $3,883.79 o $1,832.60 Suppose Olivia had deposited another $1,700 into a savings account at a second bank at the same time. The second bank also pays a nominal (or stated) interest rate of 7.8% but with quarterly compounding. Keeping everything else constant, how much money will Olivia have in her account at this bank in 11 years? O $1,836.53 O$334.34 O $3,976.37 O $232.60
Answer #1

2)Amount in 11 years = Amount deposited [1+ i]^n

            = 1700 [1+.078]^11

             =1700 [1.078]^11

             = 1700 * 2.28458

              = 3883.79

correct option is "C"

3)Quarterly rate= 7.8/4 = 1.95%

number of quarters= 11*4=44

amount in 11 years = 1700 [1+.0195]^44

              = 1700* 2.33904

                 = 3976.37

correct option is "C"